# Scaling Optimal LR Across Token Horizon

### v1

@@ Abstract
State-of-the-art LLMs are powered by scaling – scaling model size, dataset size and cluster size. It is economically infeasible to extensively tune hyperparameter for the largest runs. Instead, approximately optimal hyperparameters must be inferred or transferred from smaller experiments. Hyperparameter transfer across model sizes has been studied in Yang et al. (2022). However, hyperparameter transfer across dataset size – or token horizon – has not been studied yet. To remedy this we conduct a large scale empirical study on how optimal learning rate (LR) depends on token horizon in LLM training. We first demonstrate that the optimal LR changes significantly with token horizon – longer training necessitates smaller LR. Secondly we demonstrate the the optimal LR follows a scaling law, and that the optimal LR for longer horizons can be accurately estimated from shorter horizons via our scaling laws. We also provide a rule-of-thumb for transferring LR across token horizons with zero overhead over current practices. Lastly we provide evidence that LLama-1 used too high LR, and estimate the performance hit from this. We thus argue that hyperparameter transfer across data size is an important and overlooked component of LLM training.
@@ 1 Introduction
Figure 1: Final validation loss for different learning rates (LR) and token horizons. The dashed lines indicate our fitted curve and the stars indicate estimated optimal LR. The optimal LR decreases for as the token horizon increases.
State-of-the-art LLMs are scaled in multiple dimensions. The models are becoming increasingly large, e.g. Grok-1.5 has 314 billion (B) parameters (xAI, 2024). The clusters used to trained them are growing in size, e.g. the recently operational Memphis super-cluster contains over 100,000 H100 GPUs (Alcorn, 2024). Lastly the training datasets are growing, e.g. LLama-3 was trained on 15 Trillion (T) tokens (Dubey et al., 2024). At these scales it is infeasible to extensively tune hyperparameters. Practitioners must instead resort to hyperparameter transfer, a process where approximately optimal hyperparameters for large scale experiments are inferred from experiments at a smaller scale. The perhaps most famous work on hyperparameter transfer is muP (Yang et al., 2022) – a methodology for transferring optimal hyperparameter from a small model to a large model. While hyperparameter transfer across model size is a well studied problem, transfer across dataset size remains unstudied. This paper aims to remedy this shortcoming in the literature.
In this paper we present a large-scale study on hyperparameter transfer across dataset size. We specifically focus on learning rate (LR), an important hyperparameter which influences training stability, generalization and convergence speed. Our study is essentially a large ablation experiment where we vary LR and token horizon for a few different LLM models. To keep the scope and computational requirements manageable we focus on how LR depends on token horizon, an only present some preliminary results on its interaction on model size. Our experiments are based on standard public resources – the Megatron codebase (Shoeybi et al., 2019) and the refinedweb dataset (Penedo et al., 2023). Within this framework we have three main contributions:
@@
  • We demonstrate that 1) the optimal LR for any given architecture follows a scaling law and that 2) this allows hyperparameter transfer where the optimal LR for a long horizon is inferred from a shorter horizon. Furthermore we provide a rule-of-thumb for transferring LR across token horizons with zero overhead over current practices.
  • We provide a case-study in the optimal LR for the LLama architecture. We provide evidence that LLama-1 used a LR which is too large, highlighting hyperparameter transfer across horizons as an overlooked component of LLM training.
@@ 2 Background
LLM Scaling Laws. LLMs are typically decoder-only transformers Vaswani (2017) trained via next token prediction on web-scale text Radford et al. (2019). It has been empirically observed that the performance of LLMs scale well with model size – with the largest model showing emergent capabilities Wei et al. (2022). Kaplan et al. (2020) shows that LLM performance roughly scales as a power-law in the model size $N$. Performance here is measured by validation loss $L$, which is well known to correlate strongy with downstream metrics. Specifically they propose the following law (using constants $N_{c},\alpha_{N}$) for models trained on sufficiently large datasets:
$L(N)=(N_{c}/N)^{\alpha_{N}}$ (1)
Hoffmann et al. (2022) also showed that the performance scales well with dataset size and that the optimal performance for a given FLOPs budget is obtained by scaling model and dataset size jointly. The current paradigm thus scales data in addition to model size – e.g. Llama-1 (Touvron et al., 2023) was trained on 1.4 trillion tokens while LLama-3 used 15 trillion tokens (Dubey et al., 2024). In this paper we use notation from Kaplan et al. (2020), denoting the dataset size (i.e. number of tokens) by $D$ and the model size (i.e. parameter count) by $N$. As is common in the literature, we will fit scaling laws to empirical observations. To measure goodness-of-fit we will use the $R^{2}$ measure – its value ranges from 0 to 1, where 1 indicates a perfect fit and 0 indicates no fit at all.
Hyperparameter transfer. Hyperparameters can strongly influence the performance of LLMs, and LR is a critical hyperparameter. The muP paper Yang et al. (2022) first popularized hyperparameter transfer – the process of finding optimal hyperparameters from a small proxy experiment. Core to muP is the muP-parametrization – a way of parametrizing an LLM such that the learning rate which is optimal for a small model is also optimal for a larger model. The muP parametrization introduces some changes to the network, e.g. the attention scaling factor is changed from $\frac{1}{\sqrt{d}}$ to $\frac{1}{d}$. The original muP paper shows that LR and many other hyperparameters transfer from small models to large models under muP parametrization. This is further corroborated in Lingle (2024).
Figure 2: Final validation loss as a function of learning rate (LR) and token horizon. The dashed lines indicate our fitted curve and the stars indicate optimal LR. The optimal LR decreases monotonically for longer horizons.
@@ 3 Experiments
Experimental Setup. Our setup follows standard training of decoder-only transformer LLMs. We use the Megatron codebase Shoeybi et al. (2019), the model sizes of Table 7, and hyperparamaters roughly following GPT-3 Brown (2020) – weight decay of 0.1, gradient clipping of 1.0, and cosine learning decay schedule. The full list of hyperparameters can be viewed in Table 6 in Appendix A. We make three adjustments compared the GPT-3, mainly to prevent divergence of the model for high learning rate. 1) The GPT-3 paper does not specify the batch-size schedule, and thus not the complete warm-up schedule. Given this we use the maxima of 1% of the training steps following Muennighoff et al. (2024) and 1000. Having too short warmup stage is known to cause divergence Wortsman et al. (2023) which we want to avoid for short training runs. 2) We use qk-norm following Dehghani et al. (2023), this is known to prevent divergence without hurting validation loss Wortsman et al. (2023). 3) The original GPT-3 paper use different batch sizes for different model sizes. To avoid confounders we simply use the same batch size of 0.5m tokens all model. We will use a cosine-decay schedule and thus for each run the LR will vary with the training step. We will use the term LR to refer to the largest LR. For learning rates we will start with the base LRs from Table 7 and potentially multiply it by some factor. We use the RefinedWeb dataset (Penedo et al., 2023), a common-crawl derived dataset of roughly 600B tokens which is known to be of high quality (Penedo et al., 2024). For our experiments we will consider token horizons measured in billions of tokens, and the highest horizon we consider is 800B. In this setting the model will see each token either once or twice, and such few exposures are known to not be harmful (Muennighoff et al., 2024). Unless mentioned we will use the same seed for all runs. This implies that models with the same token horizon will see exactly the same data. For curve fitting we use least-square fitting with either a first or second degree polynomial using numpy and scipy (Harris et al., 2020).
@@
Figure 3: Final validation loss as a function of max learning rate (LR) and token horizon for four models. The dashed lines indicate our fitted curve. The optimal LR, denoted by a black star, decreases monotonically with longer horizons for all models. .
3.1 Initial Ablations
@@
In our first experiment we consider the 350m model from Table 7. We perform an ablation study where we vary the LR and token horizon and measure the final validation loss. We consider $\{25,50,100,200,400\}$ billion tokens. We will start with the base LR of $3\!\times\!10^{-4}$ from Table 7 and then multiply it by factors $\{0.25,0.5,1,2,4\}$. The LRs we consider are thus $\{7.5\!\times\!10^{-5},1.5\!\times\!10^{-4},3\!\times\!10^{-4},6\!\times\!10^{-4},1.2\!\times\!10^{-3}\}$. For each combination of LR and token horizon we train a model and record its final validation loss. To make sure we have the best possible resolution around the minima we find the minimizer $LR^{*}$ among the LR multipliers $\{0.25,0.5,1,2,4\}$, and then further add the LRs halfway between the minimizer and its two nearest neighbor – a procedure we repeat twice. From these losses we fit a curve which estimates how the final validation loss depends on the LR. For each token horizon we fit a second degree polynomial in $\log(LR)$ which provides an excellent fit. The $R^{2}$ of the fit are 0.995 or better, see Table 9 in the Appendix for exact values. The validation losses and fitted curve are shown in Figure 2. We also estimate the optimal LR by taking the minimizer of the fitted curve. From Figure 2 we make two observations: 1) the optimal LR decreases with longer token horizons, 2) the quadratic fit provides excellent agreement with experimental data.
@@
We now repeat these experiments for more model sizes. Specifically we chose the model sizes 50m, 125m, 700m and 1b from Table 7. For computational reasons we only consider shorter horizons for the larger models - the 50m and 125m models go up to 800B tokens, the 760m and 1.3B models goes up to 200B tokens while the 2.7B model only goes up to 200m tokens. For the larger models we also don’t increase the sampling around the minimizer. The results are shown in Figure 3. Two runs (model sizes 50m and 125m, 800B tokens) diverge, and we remove these. The results look similar to those of Figure 2 and we observe that the optimal LR decreases with longer token horizons across model sizes. Figure 11 in the Appendix show similar results. We thus conclude that this is a robust phenomena.
@@
Figure 4: Scaling laws for optimal LR versus token Horizon. We compare the empirically best LR (dots) to the smooth scaling law of Equation 3 with fitted constants. The $R^{2}$ of these fits are in the range 0.99 - 0.96. Across all model sizes we see that the scaling law of provides a good fit to empirical data. Note the logarithmic axes.
@@ 3.2 Scaling Laws
We have seen that longer token horizons require smaller LR. We now investigate if this insight allows us to do hyperaprameter-transfer – i.e. finding the optimal LR to a long token horizon from experiments on a shorter horizon. To do this we will fit scaling laws to our empirical investigations. Given some fixed model architecture and training recipe, let ${LR}^{*}(D)$ denote the optimal LR for some token horizon $D$. Following Kaplan et al. (2020) we will use the following functional form
@@
Here $B$ and $\beta$ are two constants which might e.g. depend on the model architecture. Taking the logarithm of both sides of Equation 3 we get
@@
We thus have a linear equation in the unknowns $\log B,\beta$, and can fit the unknown variables with least squares from our empirical data. Fitting this constant with least squares give reasonable fits – the curves for four models are shown in Figure 4. The $R^{2}$ of these fits are in the range 0.99-0.96 (see Table 8 in Appendix A for exact values), a rather high number considering the measurement variance of Section 4.4. In Figure 12 in the Appendix we also show the fits for the 1.3B and 2.7B models, but these use shorter token horizons.
To evaluate the scaling laws we cannot solely rely on $R^{2}$ as that is essentially a measure of training loss. We instead need to evaluate the the fit on some held-out data. To simulate hyperparameter transfer we thus consider fitting the constants $\log B,\beta$ on token horizons $25B,50B$ and $100B$ – and then use these constants to predict the optimal LR at larger horizons. We will use the optimal LRs we have empirically found to estimate what the correct LR is. This comparison is illustrated for the 125m model in Table 1. We see a reasonably good fit with a relative error of 10-15% – a clear improvement over not changing LR at all when going to longer horizons. We thus conclude that it is possible to perform hyperparameter transfer across token horizons with our scaling laws.
@@
Table 1: Simulated hyperparameter transfer accross token horizons for a 30m model. We use the empirically measured optimal LR at 25,50 and 100B tokens to estimate the optimal LR at longer horizons by fitting the constants of Equation 3. We find a reasonable fit when scaling up the horizon to 800B tokens, with a relative error of 10-15 %. The relative error of using the best LR at the 100B horizon for a 800B horizon is $>$ 250%.
4 Additional Experiments
@@
While we have demonstrated that the optimal LR decreases with token horizons it is natural to ask – does muP alleviate this issue? I.e. Does muP transfer across token horizons? We here investigate if that is the case. To do this we consider the 125m model from Table 7, and use muP parameterization. Note, we will not perform hyperparameter transfer across model sizes, we just use the muP parameterization which will slightly differ from the standard parameterization of transformers. We then repeat then perform ablation experiments where we vary the token horizon and the LR, using horizons $\{25,50,100\}$ billion tokens and LRs of $\{0.25,0.5,1,2,4\}$ times the base LR in Table 7. The results of this experiment are shown in Figure 5, and we can indeed see that the optimal LR decreases with token horizon. This demonstrates that the optimal LR does not transfer across token horizon with muP.
@@
Figure 5: Optimal LR vs token horizon for a tiny model using muP parameterization (Yang et al., 2022). We see that the optimal LR decreases with longer token horizon, thus suggesting that LR does not transfer across horizons even with muP.
4.2 Jointly Fitting Token Horizon and Model Size
$LR^{*}(N,D)=B(N)D^{-\beta(N)}$ (4)
$LR^{*}(N,D)=A(D)N^{-\alpha(D)}$ (5)
The constants $A,\alpha$ may depend on $D$.
The key question that arises is how the general notion of model size $N$ can be incorporated into the joint scaling law. Moreover, the scaling law formula from Eq. 5 for constant $D$ has to be representable by Eq. 4. It is anticipated to align with the latter, consisting of distinct power laws, each with specific parameters for different $N$ and $D$ values. Consequently, the objective is to identify a function that fulfills these criteria
$LR^{*}(N,D)=B(N)D^{-\beta(N)}=A(D)N^{-\alpha(D)}$ (6)
Figure 6: (a) There is a linear relationship between $\log D$ and $\log LR^{*}$ for different values of model size $N$. Both axes are in the log-scale. (b) There is a linear relationship between $\log N$ and $\log LR^{*}$ for different values of dataset size $D$. Both axes are in the log-scale.
$B(N)N^{\alpha}=A(D)D^{\beta}$ (7)
The LHS of Eq. 7 only depend on $N$, whereas the RHS only depends on $D$ so they should both equal some constant,$C$ (this step relies on our proof above that $\alpha,\beta$ $N,D$), resulting in the functional forms of $A(D),B(N)$
$A(D)=CD^{-\beta},B(N)=CN^{-\alpha}$ (8)
Plugging the functional forms of $A(D),B(N)$ we finally get the final functional form for the joint scaling law:
$LR^{*}(N,D)=CN^{-\alpha}D^{-\beta}$ (9)
Figure 7: : Fit of the scaling laws compared to the experimental results.
Subsequently, we fit parameters in Eq. 9 to describe the Optimal learning rate scaling. For comparison, similar to (Hoffmann et al., 2022), we use Huber loss (Huber, 1964), with $\delta=1\times 10^{-3}$. The optimization is performed using the BFGS algorithm. We tested for different seeds. We observe a good fit, with RMSE$=2.7\times 10^{5}$. The values are presented in Table 2. We depict the results in Figure 7.
$C$ $\alpha$ $\beta$
0.0077 0.23 0.32
Table 2: Values of the fitted coefficients. Those values assume $D$ in Billion and $N$ in Million.
4.3 A Case-study on Llama
We now consider evaluating if the LRs used by Llama-1 Touvron et al. (2023) are ”correct” according to our scaling laws. To do this we adopt the LLama-1 architecture (RMSnorm, Rope embeddings and so on, see Touvron et al. (2023) for details) and run small scaling experiments with token horizons 25B, 50B and 100B. Based upon these experiments, we find that values for Figure 8 are $B=8.29\times 10^{-4}$,$\beta=0.2964$. This results in the following prediction for optimal learning rate for different values of token horizon for this model as can be seen in Table 3. Note that we predict that the optimal LR at 1T tokes is $1.15\!\times\!10^{-4}$, whereas LLama-1 used $3\!\times\!10^{-4}$. Thus, our results suggests that the original Llama paper used an LR which was too large by a factor of $>2.5$. We argue that this demonstrates how hyperparameter transfer across token size is an overlooked component for LLM training. Notably, we also predict the optimal learning rate at $N=6700M,D=1000B$ using our scaling law from Eq. 9) and get $1.1\times 10^{-4}$.
Token Horizon Optimal LR
25B $3.4\times 10^{-4}$
50B $2.8\times 10^{-4}$
100B $2.3\times 10^{-4}$
1T $1.15\times 10^{-4}$
1T (Using Eq. 9) $1.11\times 10^{-4}$
Table 3: Optimal LR for LLama-1 for different token horizon using the fit in Figure 8. Note that the real Llama-1 used a LR of $3\!\times\!10^{-4}$ which our results suggests is significantly too large.
4.4 Uncertainty Quantification

### v2

@@ Abstract
State-of-the-art LLMs are powered by scaling – scaling model size, dataset size, and cluster size. It is economically infeasible to extensively tune hyperparameters for the largest runs. Instead, approximately optimal hyperparameters must be inferred or transferred from smaller experiments. Hyperparameter transfer across model sizes has been studied in Yang et al. (2022). However, hyperparameter transfer across dataset size – or token horizon – has not been studied yet. To remedy this we conduct a large-scale empirical study on how optimal learning rate (LR) depends on the token horizon in LLM training. We first demonstrate that the optimal LR changes significantly with token horizon – longer training necessitates smaller LR. Secondly, we demonstrate that the optimal LR follows a scaling law and that the optimal LR for longer horizons can be accurately estimated from shorter horizons via such scaling laws. We also provide a rule-of-thumb for transferring LR across token horizons with zero overhead over current practices. Lastly, we provide evidence that LLama-1 used too high LR, and argue that hyperparameter transfer across data size is an overlooked component of LLM training.
@@ 1 Introduction
Figure 1: Final validation loss of a 350 million parameter LLM for different learning rates (LR) and token horizons. The dashed lines indicate our fitted curve and the stars indicate the estimated optimal LR. The optimal LR decreases as the token horizon increases.
State-of-the-art LLMs are scaled in multiple dimensions. The models are becoming increasingly large, e.g. Grok-1.5 has 314 billion (B) parameters (xAI, 2024). The clusters used to train them are growing in size, e.g. the recently operational Memphis super-cluster contains over 100,000 H100 GPUs (Alcorn, 2024). Lastly, the training datasets are growing, e.g. LLama-3 was trained on 15 Trillion (T) tokens (Dubey et al., 2024). At these scales, it is infeasible to extensively tune hyperparameters. Practitioners must instead resort to hyperparameter transfer, a process where approximately optimal hyperparameters for large-scale experiments are inferred from experiments at a smaller scale. Perhaps the most famous work on hyperparameter transfer is muP (Yang et al., 2022) – a methodology for transferring optimal hyperparameter from a small model to a large model. While hyperparameter transfer across model size is a well-studied problem, transfer across dataset size remains understudied. This paper aims to remedy this shortcoming in the literature.
In this paper, we present a large-scale study on hyperparameter transfer across dataset sizes. We specifically focus on learning rate (LR), an important hyperparameter that influences training stability, generalization, and convergence speed. Our study is essentially a large ablation experiment where we vary LR and token horizon for a few different LLM models. We consider $>$250 training runs in total. To keep the scope and computational requirements manageable we focus on how LR depends on the token horizon, and only present some preliminary results on its interaction on model size. Our experiments are based on standard public resources – the Megatron codebase (Shoeybi et al., 2019) and the RefinedWeb dataset (Penedo et al., 2023). Our three main contributions are:
@@
  • We demonstrate that 1) the optimal LR for any given architecture follows a scaling law and 2) this allows hyperparameter transfer where the optimal LR for a long horizon is inferred from a shorter horizon. Furthermore, we provide a rule-of-thumb for transferring LR across token horizons with zero overhead over current practices.
  • We provide a case study on the optimal LR for the LLama architecture. We provide evidence that LLama-1 used an LR that is too large, highlighting hyperparameter transfer across horizons as an overlooked component of LLM training.
@@ 2 Background
[…]
Hyperparameter transfer. Hyperparameters can strongly influence the performance of LLMs, and LR is a critical hyperparameter. The muP paper (Yang et al., 2022) first popularized hyperparameter transfer – the process of finding optimal hyperparameters from a small proxy experiment. Core to muP is the muP-parametrization – a way of parametrizing an LLM such that the optimal LR for a small model is also optimal for a larger model. The muP parametrization introduces some changes to the network, e.g. the attention scaling factor is changed from $\frac{1}{\sqrt{d}}$ to $\frac{1}{d}$. The original muP paper shows that LR and many other hyperparameters transfer from small to large models when muP parametrization is used, which is further corroborated in Lingle (2024). While Yang et al. (2022) focuses on transfer between model sizes, we focus on transfer between token horizons.
Figure 2: Final validation loss as a function of learning rate (LR) and token horizon. The dashed lines indicate our fitted curve and the stars indicate optimal LR. The optimal LR decreases monotonically with longer horizons.
@@ 3 Experiments
Experimental Setup. Our experimental setup closely follows the setup of the GPT-3 paper (Brown, 2020). We consider model sizes of 50 million (m), 125m, 350m, 760m, 1.3 billion (B), and 2.7B parameters using the architectures of Table 2.1 in the GPT-3 paper. The table is replicated as Table 4 in Appendix A and also lists the LRs used in the GPT-3 paper. We use hyperparameters following GPT-3 – weight decay of 0.1, gradient clipping of 1.0, and cosine learning decay schedule. The full list of hyperparameters can be viewed in Table 3 in Appendix A. The LR will vary with the training step since we use an LR schedule, and we will use the term LR to refer to the largest LR of a run. We make three adjustments compared to the GPT-3, mainly to prevent divergence of the model for high LRs. 1) For warmup we use the maxima of 1% of the training steps following Muennighoff et al. (2024) and 1000. Having too short a warmup stage is known to cause divergence (Wortsman et al., 2023) which we want to avoid for short training runs. 2) We use qk-norm following Dehghani et al. (2023), this is known to prevent divergence without hurting validation loss (Wortsman et al., 2023). 3) The original GPT-3 paper uses different batch sizes for different model sizes. To avoid confounders we simply use the same batch size of 0.5m tokens all model. We use the RefinedWeb dataset (Penedo et al., 2023), a common-crawl derived dataset of roughly 600B tokens which is known to be of high quality (Penedo et al., 2024). Experiments are run on the Megatron codebase (Shoeybi et al., 2019). Unless mentioned we will use the same seed for all runs. For curve fitting we use least-square fitting with either a first or second-degree polynomial using Numpy and Scipy (Harris et al., 2020).
@@
Figure 3: Final validation loss as a function of max learning rate (LR) and token horizon for four models. The dashed lines indicate our fitted curve. The optimal LR, denoted by a black star, decreases monotonically with longer horizons for all models.
3.1 Ablations
@@
In our first experiment, we consider the 350m LLM model from Table 4 in Appendix A. We perform an ablation study where we vary the LR and token horizon and measure the final validation loss. We consider $\{25,50,100,200,400\}$ billion tokens. We will start with the LR of $3\!\times\!10^{-4}$ from Table 4 which is what the GPT-3 paper used (Brown, 2020). We then multiply this base LR by factors $\{0.25,0.5,1,2,4\}$. The LRs we consider are thus $\{7.5\!\times\!10^{-5},1.5\!\times\!10^{-4},3\!\times\!10^{-4},6\!\times\!10^{-4},1.2\!\times\!10^{-3}\}$. For each combination of LR and token horizon, we train a model and record its final validation loss. To make sure we have the best possible resolution around the minima we find the optimal learning rate $LR^{*}$, and then further train with the LRs halfway between $LR^{*}$ and its two nearest neighbors – a procedure we repeat twice. From these losses, we fit a curve that estimates how the final validation loss depends on the LR. For each token horizon we fit a second degree polynomial in $\log(LR)$ which provides an excellent fit. The $R^{2}$ of the fit are 0.995 or better, see Table 6 in the Appendix for exact values. The validation losses and fitted curve are shown in Figure 2. We also estimate the optimal LR by taking the minimizer of the fitted curve. From Figure 2 we make two observations: 1) the optimal LR decreases with longer token horizons, and 2) the quadratic fit provides excellent agreement with experimental data.
@@
We now repeat these experiments for more model sizes – specifically the 50m, 125m, 760m, 1.3B, and 2.7B parameter models from Table 4. For computational reasons, we only consider shorter horizons for the larger models. The 50m and 125m models go up to 800B tokens, the 760m and 1.3B models go up to 200B tokens while the 2.7B model only goes up to 100B tokens. For the larger models, we don’t increase the sampling rate around the minimizer. For each model, we consider the base LR as the one used in the GPT-3 paper (also listed in Table 4) and then multiply it by e.g. $\{0.25,0.5,1,2,4\}$. The results of these experiments are shown in Figure 3. Two runs (model sizes 50m and 125m, 800B tokens, and highest LR) diverge, and we remove these. The results look similar to those of Figure 2 and we observe that the optimal LR decreases with longer token horizons across model sizes. Figure 10 in the Appendix show similar results for the 1.3B model. We thus conclude that this is a robust phenomenon.
@@
Figure 4: Scaling laws for optimal LR versus token Horizon. We compare the empirically best LR (dots) to the smooth scaling law of Equation 1 with fitted constants. The $R^{2}$ of these fits are in the range 0.99 - 0.96. Across all model sizes, we see that the scaling law provides a good fit to the empirical data.
@@ 3.2 Scaling Laws
We have seen that longer token horizons require smaller LR. We now investigate if this insight allows us to derive scaling laws and do hyperaprameter-transfer – i.e. finding the optimal LR to a long token horizon from experiments on a shorter horizon. To do this we will fit scaling laws to our empirical results. Given some fixed model architecture and training recipe, let ${LR}^{*}(D)$ denote the optimal LR for some token horizon $D$. We will use the following functional form:
@@
Here $B$ and $\beta$ are two constants that might e.g. depend on the model architecture. Taking the logarithm of both sides of Equation 1 we get
@@
We thus have a linear equation in the unknowns $\log B,\beta$, and can fit these to our experimental results with least squares. Specifically, we use the minimizer of the quadratic fits of Figure 3 as the data points $LR^{*}$. Fitting Equation 2 gives good fits – the curves for four models are shown in Figure 4. The $R^{2}$ of these fits are in the range 0.99-0.96 (see Table 5 in Appendix A for exact values), a rather high number. In Figure 11 in the Appendix we also show the fits for the 1.3B and 2.7B models, which also show a good fit.
To evaluate the scaling laws we cannot solely rely on $R^{2}$ as that is essentially a measure of training loss. We instead need to evaluate the fit on some held-out data. To simulate hyperparameter transfer we thus consider fitting the constants $\log B,\beta$ on token horizons $25B,50B$ and $100B$ – and then use these constants to predict the optimal LR at longer horizons. We will use the optimal LR we have empirically found with quadratic fits as the correct LR. The results are illustrated for the 50m model in Table 1. We see a reasonably good fit with a relative error of 10-15% – a clear improvement over not scaling LR at all which has a relative error of $>250\%$. We thus conclude that it is possible to perform hyperparameter transfer across token horizons with our scaling laws. Table 1 is repeated for more models with similar fits in Appendix B.
@@
Table 1: Simulated hyperparameter transfer across token horizons for a 50m model. We use the empirically measured optimal LR at 25,50 and 100B tokens to estimate the optimal LR at longer horizons by fitting the constants of Equation 2. We find a reasonable fit when scaling up the horizon to 800B tokens, with a relative error of 10-15 %. The relative error of using the best LR at the 100B horizon for an 800B horizon is $>$ 250%.
@@
It is natural to ask if muP allows hyperparameter transfer across token horizons. To investigate this we consider the 50m model from Table 4, and use muP parameterization. Note, we will not perform hyperparameter transfer across model sizes, we just use the muP parameterization which slightly differs from the standard parameterization of transformers. We then run ablation experiments using token horizons $\{25,50,100\}$ billion tokens and LRs of $\{0.25,0.5,1,2,4\}$ times the base LR in Table 4. The results of this experiment are shown in Figure 5, and we can indeed see that the optimal LR decreases with a longer token horizon. We thus conclude that the optimal LR does not transfer across token horizons with muP.
@@
Figure 5: Optimal LR vs token horizon for a 50m model using muP parameterization (Yang et al., 2022). We see that the optimal LR decreases with longer token horizons, demonstrating that LR does not transfer across horizons even with muP.
3.4 Quantifying Variance